# Correctness Assessment of Code Generated by Large Language Models Using Internal Representations

### v1

@@ Abstract
Ensuring the correctness of code generated by Large Language Models (LLMs) presents a significant challenge in AI-driven software development. Existing approaches predominantly rely on black-box (closed-box) approaches that evaluate correctness post-generation, failing to utilize the rich insights embedded in the LLMs’ internal states during code generation. In this paper, we introduce Openia, a novel white-box (open-box) framework that leverages these internal representations to assess the correctness of LLM-generated code. Openia systematically analyzes the intermediate states of representative open-source LLMs specialized for code, including DeepSeek-Coder, CodeLlama, and MagicCoder, across diverse code generation benchmarks. Our empirical analysis reveals that these internal representations encode latent information, which strongly correlates with the correctness of the generated code. Building on these insights, Openia uses a white-box/open-box approach to make informed predictions about code correctness, offering significant advantages in adaptability and robustness over traditional classification-based methods and zero-shot approaches. Experimental results demonstrate that Openia consistently outperforms baseline models, achieving higher accuracy, precision, recall, and F1-Scores with up to a 2X improvement in standalone code generation and a 46% enhancement in repository-specific scenarios. By unlocking the potential of in-process signals, Openia paves the way for more proactive and efficient quality assurance mechanisms in LLM-assisted code generation.
Index Terms:
@@ 1 Introduction
The emergence of Large Language Models for Code (Code LLMs) has transformed the landscape of automated code generation, offering a promising solution to improve developer productivity and address the growing demand for software [1, 2, 3, 4]. These models, trained on vast repositories of code, are capable of generating complex functionalities, suggesting optimizations, and even automating repetitive tasks. However, as LLM-generated code becomes increasingly integrated into real-world systems, ensuring the quality and reliability of this code has become critical [5, 6, 7, 8]. Buggy/vulnerable code can lead to significant functional failures, increased maintenance costs, and severe security vulnerabilities, emphasizing the need for rigorous evaluation and assurance mechanisms [9].
Despite this, recent studies show a concerning trend that LLM-generated code frequently contains more bugs and security vulnerabilities than human-written code [10, 11, 12, 7, 13, 8, 14, 5]. This problem has led developers to find LLM-generated code often unusable, resulting in considerable time spent debugging or discarding it entirely [15, 1]. Current approaches to detect and mitigate these errors rely heavily on post-hoc analyses, such as static or dynamic code analysis and testing, which can be resource-intensive and may miss context-specific issues.
Human coding processes remain a “closed-box”, where we only see the final code, with no insights into the developer’s thought process or intermediate steps. In contrast, AI (LLMs) code generation is inherently accessible: we can examine the model’s internal states and byproducts, especially with open-source code LLMs [16, 17, 18], providing a transparent, “open-box” advantage unique to code LLMs. This visibility opens an exciting opportunity: Can we leverage this visibility to proactively detect and correct issues during the code generation process, preventing bugs and vulnerabilities before they reach developers?
@@
We hypothesize that the internal states of LLMs encapsulate meaningful signals that reflect the reliability of the code they generate. These signals, embedded in the representations formed during the generation process, may correlate with key quality attributes. Understanding and leveraging these signals could enable the proactive identification of faults, thereby improving the overall quality assurance pipeline.
In this work, we tackle the challenges and gaps in assessing the reliability of code generated by LLMs by presenting two key contributions. First, we introduce Openia, a novel framework that leverages the internal representations of code LLMs during code generation to assess the correctness of the generated code. Unlike traditional quality assurance techniques that rely heavily on post-hoc validation, Openia identifies and utilizes the signals embedded within the LLM’s internal states. By doing so, it enables earlier detection of potential bugs, reducing both the computational cost and the time required for quality assurance. This proactive approach enhances the utility of LLMs in generating high-quality code for real-world applications. Second, we conduct a systematic analysis of the internal signals of code LLMs during code generation, focusing on three representative open-source models, including DeepSeek Coder, Code Llama, and Magic Coder. These models are evaluated across diverse benchmarks, encompassing standalone code generation and repository-level generation tasks. Our analysis reveals specific patterns and features in the internal representations that align with the reliability attributes of the generated code. Specifically, our experimental results demonstrate that Openia consistently outperforms traditional classification-based and zero-shot approaches across a range of benchmarks. For independent/standalone code generation tasks, Openia achieves up to a 2X improvement in accuracy compared to zero-shot baselines and surpasses classification-based methods, such as those utilizing CodeBERT and CodeT5+, by up to 20% in accuracy. In repository-level code generation, Openia shows a remarkable, up to 46%, improvement in F1-Score, highlighting its robustness in handling complex and context-dependent coding scenarios. These findings provide valuable insights into how LLMs encode and process information related to code correctness, offering a deeper understanding of their inner workings.
@@ 2 Internal Representations in LLMs for Code Generation
Code LLMs rely on a sequence-to-sequence mapping mechanism, where input sequences (e.g., natural language descriptions or partially written code) are mapped to output sequences (e.g., complete or corrected code) [23, 19, 20, 24]. Central to this process is the internal representations, which encode the intermediate understanding of the model as it processes inputs and generates outputs. These representations provide a window into the LLM’s decision-making process, offering insights into how the model predicts the next code tokens.
@@ 2.1 The Code Generation Process
For a code LLM $\mathscr{M}$, given an input sequence $x=\{x_{1},x_{2},...,x_{n}\}$, where $x_{i}$ represents the $i$-th token in the input, $\mathscr{M}$ generates an code output sequence $c=\{c_{1},c_{2},...,c_{m}\}$ token by token. At each generation step $s$, the model predicts the next code token $c_{s}$ by leveraging both the input sequence and the tokens generated so far, denoted $c_{<s}=\{c_{1},c_{2},...,c_{s-1}\}$. The prediction process at step $s$ can be represented as:
@@ 2.2 Internal Representations
where $h_{l-1,s}$ is the activation from the preceding layer $(l-1)$. Meanwhile, $\text{context}_{s}$ represents contextual information (e.g., attention-weighted representations of other tokens), $f_{l}$ is the layer-specific transformation function (e.g., feedforward, attention, or normalization operations) applied at layer $l$. Note that the initial hidden state for token $c_{s}$ is $h_{0,s}$ is the embedding vector $e_{s}$ of $c_{s}$.
Each $h_{l,s}$ is a vector $\mathbb{R}^{d}$ where $d$ is the dimensionality of the LLM’s hidden states (e.g., 4,096 in the case of Code Llama). The sequence of activations across all layers and code tokens forms a hierarchical representation of the input and generated sequences, that could encode syntactic, semantic, and contextual information at varying levels of abstraction. Internal representations encapsulate the underlying structure and logic the model uses to generate code. By analyzing these representations, we gain insights into the model’s ”thought process“ and ”intuition” during code generation. These representations offer valuable cues about the correctness and reliability of the generated code $c$.
@@ 3.1 Task Formulation
Given a code LLM $\mathscr{M}$, an input prompt $x$, and the corresponding LLM-generated code $c$, the objective is to predict whether $c$ meets the correctness criteria as defined by the requirements of $x$. The correctness of $c$ is evaluated by verifying that it aligns with the intended behavior and satisfies the specified requirements provided in $x$.
Let $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$ represent a set of validation tests designed to assess whether $c$ adheres to the requirements in $x$. Each test $t\in\mathcal{T}$ returns an outcome, denoted as $t(c)$, which indicates whether $c$ passed or failed the test. Code $c$ is considered correct if and only if it passes all the tests in $\mathcal{T}$.
@@ 3.2 Framework Overview
Fig. 1 illustrates the overview of our framework, Openia, which leverages the internal representations of the LLM to assess the correctness of its generated code. Specifically, given a code LLM $\mathscr{M}$ and its output code sequence $c=\{c_{1},c_{2},...,c_{m}\}$, Openia extracts the all internal states produced during $\mathscr{M}$’s generation:
@@
where $h_{l,s}$ represents the hidden state at layer $l$ and token $c_{s}$. These internal representations are then fed to a probing classifier [25] that predicts whether the generated code $c$ is correct or secure.
@@
Layer Selection for Internal Representations. Modern LLMs often contain multiple transformer layers, each contributing uniquely to the model’s ability to understand and process various tasks. For example, DeepSeek-Coder-6.7B [16] and Code Llama-7B [17] comprise 32 layers, while the 34B version of Code Llama has 60 layers. Each layer captures different levels of abstraction and semantic representation; shallow layers often encode lower-level syntactic and lexical features, while deeper layers focus more on abstract and semantic representations. Jin et al.  [26] have demonstrated the alignment of shallow and deep layers with simpler and more complex tasks, respectively. Meanwhile, Song et al.  [27] highlights the varying importance of layers within an LLM. Thus, selecting the appropriate layer(s) among $L$ layers of $\mathscr{M}$ for extracting internal representations is crucial for achieving optimal performance in downstream tasks. An inappropriate choice of layer may fail to capture the necessary information, leading to suboptimal results. In this work, we conduct several experiments to systematically evaluate the impact of layer selection on the task of assessing the quality of the LLM’s generated code (Sec. 5.2.1).
@@
Token Selection for Internal Representations. In addition to layer selection, the choice of specific tokens among the input tokens in $x$ or generated code tokens in $c$ to extract the internal representations is also very important. Several works [28, 29, 30] have highlighted the effectiveness of probing the final token of the generated answer to assess the answer’s correctness. This token typically encapsulates the accumulated context and semantics of the entire generated sequence. Therefore, the hidden states of this token can provide valuable insights for determining whether the answer is correct or not. Meanwhile, some other approaches [31, 32] focus on the final token of the prompt, just before the response is generated. The reasoning behind this choice is that this token often reflects the model’s transition from input comprehension to output generation, and its hidden states can provide valuable signals into how well the model has understood the task requirement and its readiness to generate the output. In this work, we systematically investigate the impact of different token selection strategies on assessing the correctness of the generated code. Our goal is to understand how token-level choices affect the effectiveness of Openia and to explore the relationship between token-specific hidden states and the correctness of the generated code (Sec. 5.2.2).
@@ 4 Evaluation Design
  • RQ1. Performance Analysis: To what extent, can Openia with the internal representations of the LLM help identify the correctness of its generated code? How accurately does Openia compare to the baseline approaches?
  • RQ2. Layer and Token Selection Analysis: How do the representations at various token positions and layer depths impact Openia’s performance? Is there any correlation between the complexity of the tasks and the impact of layer and token selection?
@@ 4.1 Dataset Construction
Openia constructs the dataset using four popular benchmarks, HumanEval [33], MBPP [34], and DevEval [35]. In particular, HumanEval and MBPP contain 164 and 500 stand-alone programming problems, respectively, while DevEval consists of 1,825 tasks of repository-level code generation. For each task in these benchmarks, we employed each studied model to generate 10 candidate solutions. We then evaluate the correctness of the generated solutions. To access correctness, we utilized the test cases provided by the corresponding benchmarks. Table I summarizes the number of functional correct and incorrect solutions generated by each model for benchmarks. The full set of the generated code, the corresponding internal representations from the studied code LLMs, and labels could be found on our website [36].
TABLE I: The numbers of correct (#Cor.) and incorrect (#Inc.) solutions of benchmarks
@@ 4.2 Studied Code LLMs
In this work, experiments are conducted on three representative open-source LLMs specialized for code, including DeepSeek-Coder [16], CodeLlama [17], and MagicCoder [18]. These models were chosen due to their widespread adoption and strong performance in various code-related tasks, making them ideal candidates for analyzing how internal LLM signals correlate with the reliability of generated code. Each model represents a state-of-the-art approach, utilizing architectures and training strategies specifically designed to handle the complexities of programming languages, such as syntax, semantics, and logical consistency. To accommodate hardware resource limitations, we focus on versions of these models with parameter sizes not exceeding 7B. Despite their smaller sizes, these models retain robust capabilities for code generation, ensuring that the results of our study remain relevant to practical applications. This selection also enables a fair comparison by standardizing computational resources across experiments. For consistency and reproducibility, Openia employs their base models with pre-trained weights, which are officially published on HuggingFace. Importantly, we do not perform any fine-tuning on the models to ensure the evaluation reflects their out-of-the-box capabilities.

### v2

# definitionDefinition

@@ Abstract
Ensuring the correctness of code generated by Large Language Models (LLMs) presents a significant challenge in AI-driven software development. Existing methods predominantly rely on black-box (closed-box) approaches that evaluate correctness post-generation failing to utilize the rich insights embedded in the LLMs’ internal states during code generation. This limitation leads to delayed error detection, increased debugging costs, and reduced reliability in deployed AI-assisted coding workflows. In this paper, we introduce Openia, a novel white-box (open-box) framework that leverages these internal representations to assess the correctness of LLM-generated code. By systematically analyzing the intermediate states of representative open-source code LLMs, including DeepSeek-Coder, Code Llama, and Magicoder, across diverse code generation benchmarks, we found that these internal representations encode latent information, which strongly correlates with the correctness of the generated code. Building on these insights, Openia uses a white-box/open-box approach to make informed predictions about code correctness, offering significant advantages in adaptability and robustness over traditional blackbox methods and zero-shot approaches. Our results show that Openia consistently outperforms baseline models, achieving higher accuracy, precision, recall, and F1-Scores with up to a 2X improvement in standalone code generation and a 3X enhancement in repository-specific scenarios. By unlocking the potential of in-process signals, Openia paves the way for more proactive and efficient quality assurance mechanisms in LLM-assisted code generation.
keywords:
@@ 1 Introduction
The emergence of Large Language Models for Code (Code LLMs) has transformed the landscape of automated code generation, offering a promising solution to improve developer productivity and address the growing demand for software [1, 2, 3, 4]. These models, trained on vast repositories of code, are capable of automating repetitive tasks, suggesting optimizations, and even generating complex functionalities. As LLM-generated code becomes increasingly integrated into real-world systems, ensuring the quality and reliability of this code has become critical [5, 6, 7, 8]. Low-quality code can lead to significant functional failures, severe security vulnerabilities, and increased maintenance costs, emphasizing the need for rigorous evaluation and quality assurance mechanisms [9].
However, recent studies show a concerning trend that LLM-generated code frequently contains more bugs and security vulnerabilities than human-written code [7, 8, 10, 11, 12, 13, 14, 5]. This problem has led developers to find LLM-generated code often unusable, resulting in considerable time spent debugging or discarding it entirely [1, 15]. To detect and mitigate these errors, current approaches rely on post-hoc methods, such as static or dynamic code analysis and testing, which can be resource-intensive and may miss context-specific issues.
Unlike post-hoc approaches analyzing the final code only after code generation is complete, an in-process strategy takes advantage of the internal states capturing the “thought process” of code LLMs during code generation, enabling early detection of potential issues. This is particularly promising for code LLMs, which offer a unique “open-box” advantage to access code LLMs’ internal states, especially with open-source code LLMs [16, 17, 18]. Unlike human coding, which remains a “closed-box” process where only the final code is visible, and the developer’s intermediate reasoning and thought process are inaccessible, code LLMs generate a wealth of internal representations and byproducts that can be examined in real-time. By leveraging these signals, we can move beyond passive validation and instead, proactively assess correctness as code is generated. This opens an exciting opportunity: Can we leverage this visibility to proactively detect during the code generation process, preventing bugs before they reach developers?
@@
We hypothesize that the internal states of LLMs encapsulate meaningful signals that reflect the correctness of the code they generate. These signals, embedded in the representations formed during the generation process, may correlate with key quality attributes. Understanding and leveraging these signals could enable the proactive identification of faults, thereby improving the overall quality assurance pipeline.
In this work, we tackle the challenges and gaps in assessing the reliability of code generated by LLMs by presenting two key contributions. First, we introduce Openia, a novel framework that leverages the internal representations of code LLMs during code generation to assess the correctness of the generated code. Unlike traditional quality assurance techniques that rely heavily on post-hoc validation, Openia identifies and utilizes the signals embedded within the LLM’s internal states. By doing so, it could enable earlier detection of potential bugs, reducing both the computational cost and the time required for quality assurance. Second, we conduct a systematic analysis of the internal signals of code LLMs during code generation, focusing on three representative open-source models, including DeepSeek-Coder, Code Llama, and Magicoder. These models are evaluated across diverse benchmarks, encompassing standalone code generation and repository-level generation tasks. Our analysis reveals specific patterns and features in the internal representations that align with the reliability attributes of the generated code.
Specifically, our experimental results demonstrate that Openia consistently outperforms traditional classification-based and zero-shot approaches across a range of benchmarks. For independent/standalone code generation tasks, Openia achieves up to a 2X improvement in accuracy compared to zero-shot baselines and surpasses classification-based methods, such as those utilizing CodeBERT and CodeT5+, by up to 20% in accuracy. In repository-level code generation, Openia shows a remarkable, up to 3X, improvement in F1-Score, highlighting its robustness in handling complex and context-dependent coding scenarios. These findings provide valuable insights into how LLMs encode and process information related to code correctness, offering a deeper understanding of their inner workings.
@@ 2 Internal Representations in LLMs for Code Generation
Code LLMs rely on a sequence-to-sequence mapping mechanism, where input sequences (e.g., natural language descriptions or partially written code) are mapped to output sequences (e.g., complete or corrected code) [23, 19, 20, 24]. Central to this process is the internal representations, which encode the intermediate understanding of the model as it processes inputs and generates outputs. These representations provide a window into the code LLM’s decision-making process, offering insights into how the model predicts the next code tokens.
@@ 2.1 The Code Generation Process
For a code LLM $\mathscr{M}$, given an input sequence $x=\{x_{1},x_{2},...,x_{n}\}$, where $x_{i}$ represents the $i$-th token in the input, $\mathscr{M}$ generates a code output sequence $c=\{c_{1},c_{2},...,c_{m}\}$ token by token. At each generation step $s$, the model predicts the next code token $c_{s}$ by leveraging both the input sequence and the tokens generated so far, denoted $c_{<s>}=\{c_{1},c_{2},...,c_{s-1}\}$. The prediction process at step $s$ can be represented as:
@@ 2.2 Internal Representations
[…]
Each $h_{l,s}$ is a vector $\mathbb{R}^{d}$ where $d$ is the dimensionality of the LLM’s hidden states (e.g., 4,096 in the case of Code Llama). The sequence of activations across all layers and code tokens forms a hierarchical representation of the input and generated sequences that could encode syntactic, semantic, and contextual information at varying levels of abstraction. Internal representations encapsulate the underlying structure and logic the model uses to generate code. By analyzing these representations, we gain insights into the model’s “thought process” and “intuition” during code generation. These representations offer valuable cues about the correctness and reliability of the generated code $c$.
@@ 3.1 Task Formulation
Given a code LLM $\mathscr{M}$, an input prompt $x$, and the corresponding LLM-generated code $c$, the objective is to predict whether $c$ meets the correctness criteria as defined by the requirements provided in $x$. Specifically, the correctness of $c$ is evaluated by verifying whether $c$’s behaviors align with the intended behaviors specified in $x$.
Let $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$ represent a set of validation tests designed to assess whether $c$ adheres to the requirements in $x$. Each test $t\in\mathcal{T}$ returns an outcome, denoted as $t(c)$, which indicates whether $c$ passed or failed the test. Code $c$ is considered correct if and only if it passes all the tests.
@@ 3.2 Framework Overview
Fig. 1 illustrates the overview of our framework, which leverages the internal representations of code LLMs during code generation to assess the correctness of its generated code. Specifically, given a code LLM $\mathscr{M}$ and its output code sequence $c=\{c_{1},c_{2},\dots,c_{m}\}$, Openia extracts the all internal states produced during the decoding process of $\mathscr{M}$:
@@
where $h_{l,s}$ represents the hidden state at layer $l$ corresponding to token $c_{s}$. These internal representations are then fed to a probing classifier [25], which determines whether the generated code $c$ is correct.
(a) Prompt provided to the model
(b) Response generated by DeepSeek Coder-6.7B for the prompt in Fig. 2(a).
Figure 2: An example of independent-unit code generation for a task in the HumanEval benchmark
(a) Prompt provided to the model
(b) Response generated by DeepSeek Coder-6.7B for the prompt in Fig. 3(a).
Figure 3: An example of repo-level code generation for a task in DevEval benchmark
Fig. 2 and Fig. 3 show the examples of prompts and corresponding responses generated by DeepSeek Coder-6.7B for the independent-unit and repository-level code generation tasks, respectively. Our goal is to utilize the model’s internal representations produced during the generation process to predict the correctness of the generated code (Fig. 2(b) and Fig. 3(b)). These internal representations refer to the hidden states associated with each generated token across different layers of the model. For instance, during generation, each token (e.g., keywords like for, if or identifiers like balance in Fig. 2(b)) is represented by hidden states across all 32 transformer layers of DeepSeek Coder-6.7B. Suppose Openia chooses the final code token, i.e., False in Fig. 2(b), as the focus of analysis. From this token, Openia extracts the corresponding hidden states at a particular layer, e.g., the last ($32^{nd}$) layer, which captures high-level semantic features that reflect the model’s overall reasoning about the function. These representations are then passed to a probing classifier to predict the likelihood that the generated function is correct. To obtain informative signals for code correctness prediction, Openia aims to select appropriate tokens and extract their corresponding internal representations from a suitable layer within the model.
@@
Layer Selection for Internal Representations. Modern (code) LLMs often contain multiple transformer layers, each contributing uniquely to the model’s ability to understand and process various tasks. For example, DeepSeek-Coder-6.7B [16] and Code Llama-7B [17] comprise 32 layers, while the 34B version of Code Llama consists of 60 layers. Each layer captures different levels of abstraction and semantic representation; shallow layers often encode lower-level syntactic and lexical features, while deeper layers focus more on abstract and semantic representations. Jin et al. [26] have demonstrated the alignment of shallow and deep layers with simpler and more complex tasks, respectively. Meanwhile, Song et al. [27] highlights the varying importance of layers within an LLM. These studies emphasize that not all layers contribute equally to a given task. Thus, selecting the appropriate layer(s) among $L$ layers of $\mathscr{M}$ for extracting internal representations is crucial for achieving optimal performance in downstream tasks. An inappropriate choice of layer may fail to capture the necessary information, leading to suboptimal results. In this work, we conduct several experiments to systematically evaluate the impact of layer selection on the task of assessing the correctness of LLM-generated code (Sec. 5.2.1).
@@
Token Selection for Internal Representations. In addition to layer selection, the choice of specific tokens from which internal representations are extracted is another critical factor. Valuable signals for correctness assessment can reside in several specific tokens within the generated output or its preceding input. Several works [28, 29, 30] have highlighted the effectiveness of probing the hidden states of the final token in the generated answer to assess the answer’s correctness. This token typically encapsulates the accumulated context and semantics of the entire generated sequence. Therefore, the hidden states of this token can provide meaningful insights for determining whether the answer is correct or not. Meanwhile, some other approaches [31, 32] focus on the final token of the prompt, just before the response is generated. The reasoning behind this choice is that this token often reflects the model’s transition from input comprehension to output generation, making its hidden states indicative of how well the model has understood the task requirement and its readiness to generate the output. In this work, we also systematically investigate the impact of different token selection strategies on assessing the correctness of the generated code. Our goal is to understand how token-level choices affect the effectiveness of Openia and to explore the relationship between token-specific hidden states and the correctness of the generated code (Sec. 5.2.2).
In summary, our approach, Openia, investigates how internal representations of LLMs can be effectively utilized to assess the correctness of generated code. Openia employs a representation probing classifier, which leverages these internal representations to predict whether the generated code is correct. By systematically analyzing layer selection and token selection, we aim to identify the most informative internal representations that contribute to robust and accurate correctness assessment. Through empirical evaluations, we demonstrate how different layers and token positions impact the effectiveness of correctness assessment, providing insights into optimizing internal representation selection for improved downstream task performance.
@@ 4 Evaluation Design
RQ1. Performance Analysis: To what extent can Openia leverage the internal representations of an LLM to assess the correctness of its generated code? How accurately does Openia compare to the baseline approaches?
RQ2. Layer and Token Selection Analysis: How do the representations extracted from different token positions and layer depths impact Openia’s performance?
@@ 4.1 Dataset Construction
Openia constructs the dataset using three popular benchmarks, HumanEval [33], MBPP [34], and DevEval [35]. In particular, HumanEval and MBPP contain 164 and 500 standalone programming problems, respectively, while DevEval consists of 1,825 tasks of repository-level code generation. For each task in these benchmarks, we employed each studied model to generate 10 candidate solutions. The correctness of these generated solutions is then evaluated using the test cases provided by the respective benchmarks. Table 1 summarizes the number of correct and incorrect solutions generated by each model for the given benchmarks. The complete dataset, including all generated code, the corresponding internal representations extracted from the studied code LLMs, and their correctness labels could be found on our website [36].
Table 1: The numbers of correct (#Cor.) and incorrect (#Inc.) solutions of benchmarks
@@ 4.2 Studied Code LLMs
To ensure a fair, reproducible, and practical evaluation, we carefully selected LLMs specialized for code based on the following criteria. First, the models must be open-source to enable access to their internal representations during inference. Second, the models should be trained or fine-tuned on large-scale code corpora, as code-specialized LLMs are better suited for programming tasks and widely used in practice. Third, we limited the model size to 13B parameters to accommodate hardware resource constraints. These criteria were designed to balance model performance, accessibility, and practical usability.
Following these criteria, we selected five representative models: DeepSeek Coder-1.3B [16], DeepSeek Coder-6.7B [16], Code Llama-7B [17], Code Llama-13B [17], and Magicoder-7B [18]. These models are all open-source, widely adopted in both the research and industry, and have demonstrated strong performance in various code generation tasks. Each model employs advanced architectures and training methodologies designed to capture the complexities of programming languages, including syntax, semantics, and logical consistency. For consistency across experiments, we use their instruction-tuned variants with officially released pre-trained weights from HuggingFace and do not perform any additional fine-tuning. This setup allows us to assess the models’ out-of-the-box generalization capabilities and supports a controlled comparison of how internal representations relate to code correctness.